OpenStudy (anonymous):

-d over 3 plus 10 = -11 solve for -d

OpenStudy (jaredstone4):

So you have (-d/3)+10 = -11 Subtract 10 from both sides to get (-d/3) = -21 If you cross multiply you get (-21 * 3) = -d -d = -63

OpenStudy (anonymous):

luetzj 0 Solution for -d = -63 Solution for just d is 63 The negatives would cancel out if you're solving for d.